Assume that it takes John Connor 8 labor units to produce a tomato and 4 to produce a squash in his garden. It takes Sarah Connor 15 labor units to produce a tomato and 3 to produce a squash in her garden.

a. Which Connor has an absolute advantage in the production of tomatoes?

b. Which Connor has an absolute advantage in the production of squash?

c. Which Connor has a comparative advantage in the production of tomatoes?
Answer #1

a) John Connor has an absolute advantage in the production of tomatoes as he needs less labour units than Sarah Connor to produce a tomato in his garden.

b) Sarah Connor has an absolute advantage in the production of squash as she needs takes labour units to produce a squash in her garden than taken by John Connor.

c) John Connor's opportunity cost of producing a tomato in terms of squash= 8/4= 2 squash per Tomato

Sarah Connor's opportunity cost of Producing a tomato in terms of squash= 15/3= 5 squash per Tomato

The opportunity cost of producing a tomato in terms of squash is less for John Connor than for Sarah Connor. Hence, John Connor has a comparative advantage in production of tomatoes.